A charge particle X (-60μC) is on the x axis at the point (2,0) and a particle Y( - 100μC) is on
the x axis at the point (5,0). Calculate
i) The electric field strength and electric potential at the point (4,0). (10 marks)
ii) The coordinates of the point on the x axis where the electric field strength will be zero.
1
Expert's answer
2020-04-27T10:15:07-0400

i)


"V=\frac{kq_x}{r_x}+\frac{kq_y}{r_y}=-(9\cdot10^9)\left(\frac{60\cdot10^{-6}}{2}+\frac{100\cdot10^{-6}}{1}\right)"

"V=-1.17\cdot10^{6}V"

"E=\frac{kq_x}{r_x^2}-\frac{kq_y}{r_y^2}=-(9\cdot10^9)\left(\frac{60\cdot10^{-6}}{2^2}-\frac{100\cdot10^{-6}}{1^2}\right)"

"E=7.65\cdot10^{5}\frac{V}{m}"

ii)


"\frac{kq_x}{r_x^2}-\frac{kq_y}{r_y^2}=0\to\left(\frac{60\cdot10^{-6}}{(X-2)^2}-\frac{100\cdot10^{-6}}{(5-X)^2}\right)=0"

"X=3.3\ m, Y=0\ m"